# Finite Neural Networks as Mixtures of Gaussian Processes: From Provable Error Bounds to Prior Selection

### v1

@@ A.2 Proofs Section 5
${\mathbb{W}}_{2}^{2}({\Delta}_{\mathcal{C}}\#\mathcal{N}({{\bar{0}}},\Sigma),\mathcal{N}({{\bar{0}}},\Sigma))$ $=\sum_{l=1}^{n}\Big{(}\prod_{\begin{subarray}{c}k=1\\ k\neq l\end{subarray}}^{n}\sum_{i_{k}=1}^{|{\mathcal{C}}^{k}|}{\mathbb{P}_{i_{1}\cdot i_{2}\ldots\cdot i_{n}}^{{(k)}}}\Big{)}\Big{(}\sum_{i_{l}=1}^{|{\mathcal{C}}^{l}|}{\mathbb{P}_{i_{1}\cdot i_{2}\ldots\cdot i_{n}}^{{(l)}}}{{\mathbb{E}}_{i_{1}\cdot i_{2}\ldots\cdot i_{n}}^{{(l)}}}\Big{)}$
@@
$=\sum_{l=1}^{n}\Big{(}\prod_{\begin{subarray}{c}k=1\\ k\neq l\end{subarray}}^{n}1)\Big{(}{{{\lambda}}^{(l)}}{\mathbb{W}}_{2}^{2}({\Delta}_{{\bar{{\mathcal{C}}}^{{(l)}}}}\#\mathcal{N}(0,1),\mathcal{N}(0,1))\Big{)}$
$=\sum_{l=1}^{n}{{{\lambda}}^{(l)}}{\mathbb{W}}_{2}^{2}({\Delta}_{{\bar{{\mathcal{C}}}^{{(l)}}}}\#\mathcal{N}(0,1),\mathcal{N}(0,1)).$
To conclude the proof, note that for non-zero mean, we can follow the same procedure but apply substitution $q={{T}}({{x}}-{m})$ and define $\bar{{{c}}}={{T}}({{c}}-{m})$.
A.3 Proofs Section 6
Lemma 14: Bounds on the Spectral Norm of Random Matrices with Correlated Elements
Proof  We here generalize the result presented in the main text, and prove that for both $\rho=2$ and $\rho=1$:
${\mathbb{E}}_{{{W}}\sim p_{{W}}}[\|{{W}}\|^{\rho}]^{\frac{1}{\rho}}$ $\leq\Big{(}\sum_{i,j}{\mathbb{E}}_{\bar{{{W}}}\sim\bar{p}_{{W}}}[({\bar{{{W}}}^{{(i,j)}}})^{2}]\Big{)}^{\frac{1}{2}}+\|{{M}}\|.$ (30)
Using the triangle inequality for the spectral norm, we have
${\mathbb{E}}_{{{W}}\sim p_{{W}}}[\|{{W}}\|]={\mathbb{E}}_{\bar{{{W}}}\sim\bar{p}_{{W}}}[\|\bar{{{W}}}+{{M}}\|]\leq{\mathbb{E}}_{\bar{{{W}}}\sim\bar{p}_{{W}}}[\|\bar{{{W}}}\|]+\|{{M}}\|$
Using Minkowski’s inequality for the spectral norm, we get
${\mathbb{E}}_{{{W}}\sim p_{{W}}}[\|{{W}}\|^{2}]^{\frac{1}{2}}$ $={\mathbb{E}}_{\bar{{{W}}}\sim\bar{p}_{{W}}}[\|\bar{{{W}}}+{{M}}\|^{2}]^{\frac{1}{2}}$
$\leq{\mathbb{E}}_{\bar{{{W}}}\sim\bar{p}_{{W}}}[\|\bar{{{W}}}\|^{2}]^{\frac{1}{2}}+\|{{M}}\|$
For any matrix $\bar{{{W}}}\in{\mathbb{R}}^{n\times m}$, we have $\|\bar{{{W}}}\|\leq\|\bar{{{W}}}\|_{\mathcal{F}}$, where $\|\cdot\|_{\mathcal{F}}$ denotes the Frobenius defined by $\|\bar{{{W}}}\|_{\mathcal{F}}^{2}=\sum_{i,j}({\bar{{{W}}}^{{(i,j)}}})^{2}$. Therefore,
${\mathbb{E}}_{\bar{{{W}}}\sim\bar{p}_{{W}}}[\|\bar{{{W}}}\|^{2}]^{\frac{1}{2}}\leq{\mathbb{E}}_{\bar{{{W}}}\sim\bar{p}_{{W}}}[\|\bar{{{W}}}\|_{\mathcal{F}}^{2}]^{\frac{1}{2}}\leq\Big{(}\sum_{i,j}{\mathbb{E}}_{\bar{{{W}}}\sim\bar{p}_{{W}}}[({\bar{{{W}}}^{{(i,j)}}})^{2}]\Big{)}^{\frac{1}{2}}.$
This concludes the proof for $\rho=2$. For $\rho=1$, we additionally apply Jensen’s inequality:
${\mathbb{E}}_{\bar{{{W}}}\sim\bar{p}_{{{W}}}}[\|{{W}}\|]\leq{\mathbb{E}}_{\bar{{{W}}}\sim\bar{p}_{{{W}}}}[\|{{W}}\|_{\mathcal{F}}]\leq{\mathbb{E}}_{\bar{{{W}}}\sim\bar{p}_{{W}}}\Big{[}\Big{(}\sum_{i,j}({\bar{{{W}}}^{{(i,j)}}})^{2}\Big{)}^{\frac{1}{2}}\Big{]}\leq\Big{(}\sum_{i,j}{\mathbb{E}}_{\bar{{{W}}}\sim\bar{p}_{{W}}}[({\bar{{{W}}}^{{(i,j)}}})^{2}]\Big{)}^{\frac{1}{2}}.$
Results in Table 11: Wasserstein Distance under Push-Forward Operations in SNNs
Proposition 25 (Wasserstein Distance under Stochastic Linear Operations)
Let
$L_{{{w}}}:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{m}$ be the affine operation $L_{{w}}({{z}})={{W}}{{z}}+{{b}}$ with parameters ${{W}}\in{\mathbb{R}}^{n\times m}$ and ${{b}}\in{\mathbb{R}}^{m}$. Further, assume a distribution $p_{{w}}\in{\mathcal{P}}({\mathbb{R}}^{(n+1)\times m})$ over ${{w}}=({{W}},{{b}})$ such that for any ${{z}}\in{\mathbb{R}}^{n}$, the random variable $L_{{w}}({{z}})$ follows the probability distribution $p_{L_{{w}}({{z}})}=L_{{{w}}}({{z}})\#p_{{w}}$. Then, for $\rho\in\{1,2\}$ and probability distributions $p,q\in{\mathcal{P}}_{\rho}({\mathbb{R}}^{n})$ it holds that
${\mathbb{W}}_{\rho}({\mathbb{E}}_{{{z}}\sim p}[p_{L_{{w}}({{z}})}],{\mathbb{E}}_{\tilde{{{z}}}\sim q}[p_{L_{{w}}(\tilde{{{z}}})}])\leq{\mathbb{E}}_{({{W}},{{b}})\sim p_{w}}[\|{{W}}\|^{\rho}]^{\frac{1}{\rho}}{\mathbb{W}}_{\rho}(p,q).$ (31)
[…]
${\mathbb{W}}^{\rho}_{\rho}({\mathbb{E}}_{{{z}}\sim p}[p_{L_{{w}}({{z}})}],{\mathbb{E}}_{\tilde{{{z}}}\sim q}[p_{L_{{w}}(\tilde{{{z}}})}])$ $=\inf_{\gamma\in\Gamma(p_{{w}},p,q)}{\mathbb{E}}_{(({{W}},{{b}}),{{z}},\tilde{{{z}}})\sim\gamma}[\|{{W}}{{z}}+{{b}}-{{W}}\tilde{{{z}}}+{{b}}\|^{\rho}]$
$=\inf_{\gamma\in\Gamma(p_{{w}},p,q)}{\mathbb{E}}_{(({{W}},{{b}}),{{z}},\tilde{{{z}}})\sim\gamma}[\|{{W}}{{z}}-{{W}}\tilde{{{z}}}\|^{\rho}]$
$\leq\inf_{\gamma\in\Gamma(p_{{w}},p,q)}{\mathbb{E}}_{(({{W}},{{b}}),{{z}},\tilde{{{z}}})\sim\gamma}[\|{{W}}\|^{\rho}\|{{z}}-\tilde{{{z}}}\|^{\rho}].$
Therefore, considering that $p_{{w}}\times\Gamma(p,q)\subset\Gamma(p_{{w}},p,q)$, we obtain:
$\inf_{\gamma\in\Gamma(p_{{w}},p,q)}{\mathbb{E}}_{(({{W}},{{b}}),{{z}},\tilde{{{z}}})\sim\gamma}[\|{{W}}\|^{\rho}\|{{z}}-\tilde{{{z}}}\|^{\rho}]$ $\leq{\mathbb{E}}_{({{W}},{{b}})\sim p_{{w}}}[\|{{W}}\|^{\rho}]\inf_{\gamma\in\Gamma(p,q)}{\mathbb{E}}_{({{z}},\tilde{{{z}}})\sim\gamma}[\|{{z}}-\tilde{{{z}}}\|^{\rho}]$
$={\mathbb{E}}_{{{W}}\sim p_{{w}}}[\|{{W}}\|^{\rho}\|]{\mathbb{W}}_{\rho}^{\rho}(p,q).$

### v2

@@ B.3 Proofs of the Results in Section 7
$\underset{\begin{subarray}{c}\{{\mathcal{C}}^{1},\ldots,{\mathcal{C}}^{n}\}\subset{\mathbb{R}},\\ \sum_{i=1}^{n}|{\mathcal{C}}^{j}|=N\end{subarray}}{\mathrm{argmin}}{\mathbb{W}}_{2}\left({\Delta}_{\mathrm{Post}({\mathcal{C}}^{1}\times\ldots\times{\mathcal{C}}^{n},{{T}}^{-1})+\{{m}\}}\#\mathcal{N}({m},\Sigma),\mathcal{N}({m},\Sigma)\right)$
@@
$\qquad=\underset{\begin{subarray}{c}\{{\mathcal{C}}^{1},\ldots,{\mathcal{C}}^{n}\}\subset{\mathbb{R}},\\ \sum_{i=1}^{n}|{\mathcal{C}}^{j}|=N\end{subarray}}{\mathrm{argmin}}\sum_{j=1}^{n}{{{\lambda}}^{(j)}}{\mathbb{W}}^{2}_{2}\left({\Delta}_{{\mathcal{C}}^{j}}\#\mathcal{N}(0,1),\mathcal{N}(0,1)\right),$
$\qquad=\underset{\begin{subarray}{c}\{{\mathcal{C}}^{1},\ldots,{\mathcal{C}}^{n}\}\subset{\mathbb{R}},\\ {\mathcal{C}}^{j}=N_{j}^{*},\forall j\in\{1,\ldots,n\}\end{subarray}}{\mathrm{argmin}}\sum_{j=1}^{n}{{{\lambda}}^{(j)}}{\mathbb{W}}^{2}_{2}\left({\Delta}_{{\mathcal{C}}^{j}}\#\mathcal{N}(0,1),\mathcal{N}(0,1)\right),$
$\qquad=\left\{\underset{{\mathcal{C}}^{j}\subset{\mathbb{R}},{\mathcal{C}}^{j}=N_{j}^{*}}{\mathrm{argmin}}{\mathbb{W}}^{2}_{2}\left({\Delta}_{{\mathcal{C}}^{j}}\#\mathcal{N}(0,1),\mathcal{N}(0,1)\right)\right\}_{j=1}^{n},$
where in step 2 we used the optimal grid-configuration $\{N^{*}_{1},\ldots,N^{*}_{n}\}$ as defined in Eqn(31), and in step 3, we use the independence of objective terms across dimensions.
Appendix C Expected Lipschitz Constant for Stochastic Affine Operations
Here, we show how to compute the expected value of a Lipschitz constant ${\mathbb{E}}_{{{w}}_{k}\sim p_{{{w}}_{k}}}[{\mathcal{L}}_{L^{{{w}}_{k}}}^{2}]^{\frac{1}{2}}$ appearing in Theorem 17, in the case where $L^{{{w}}_{k}}$ represents either a fully-connected or convolutional layer, and where $\rho\in\{1,2\}$.
Fully-connected layers.
For a fully-connected layer of the form $L^{{w}}({{z}})={{W}}{{z}}+{{b}}$, the Lipschitz constant is ${\mathcal{L}}_{L^{{{w}}}}=\|{{W}}\|$, where $\|\cdot\|$ denotes the spectral norm. When the matrix W{{W}} is distributed according to $p_{{W}}\in{\mathcal{P}}_{\rho}({\mathbb{R}}^{m\times n})$, computing ${\mathbb{E}}[\|{{W}}\|^{\rho}]^{1/\rho}$ is generally intractable. Instead of using the spectral norm, we therefore consider the Frobenius norm, which yields a valid Lipschitz constant: while it is not the smallest such constant, it provides a tractable upper bound. Specifically, that
${\mathbb{E}}_{{{W}}\sim p_{{W}}}[\|{{W}}\|^{\rho}]^{1/\rho}\leq{\mathbb{E}}_{{{W}}\sim p_{{W}}}\left[\|{{W}}\|_{\mathcal{F}}^{\rho}\right]^{1/\rho}.$
This bound can be tightened by centering the distribution of W{{W}} around its mean ${{M}}={\mathbb{E}}_{{{W}}\sim p_{{W}}}[{{W}}]$ and by applying Minkowski’s inequality:
${\mathbb{E}}_{{{W}}\sim p_{{W}}}[\|{{W}}\|^{\rho}]^{1/\rho}\leq{\mathbb{E}}_{{{W}}\sim p_{{W}}}[\|{{W}}-{{M}}\|^{\rho}]^{1/\rho}+\|{{M}}\|\leq{\mathbb{E}}_{{{W}}\sim p_{{W}}}\left[\|{{W}}-{{M}}\|_{\mathcal{F}}^{\rho}\right]^{1/\rho}+\|{{M}}\|.$
Here, we again use that the spectral norm is upper bounded by the Frobenius norm.
For $\rho=2$, the upper bound simplifies in terms of second moments of the entries of W{{W}}. For $\rho=1$, we additionally apply Jensen’s inequality to ensure tractability:
${\mathbb{E}}_{{{W}}\sim p_{{W}}}\left[\|{{W}}-{{M}}\|_{\mathcal{F}}\right]\leq{\mathbb{E}}_{{{W}}\sim p_{{W}}}\left[\|{{W}}-{{M}}\|_{\mathcal{F}}^{2}\right]^{1/2}.$
Convolutional layers.
If $L^{{{w}}_{k}}$ corresponds to a 2D convolutional operation with kernel weight matrix W{{W}}, then the Lipschitz constant is given by ${\mathcal{L}}_{L^{{w}}}=\sqrt{m}\|{{W}}\|$, where $m$ denotes the maximum number of times any input element contributes to the output (e.g., the number of spatial positions a kernel covers) Gouk et al. (2021). For example, with unit stride and padding of one, $m$ equals the product of the kernel height and width. We apply the same bounding procedure as in the fully-connected setting, now including the multiplicative $\sqrt{m}$ correction factor.
Appendix D Alternative Signature Placement: Cross Scheme
As discussed in Remark 24, we constrained signature locations to axis-aligned grids in order to obtain formal error bounds (Corollary 11). When formal bounds are not required, we may instead place signature locations on a cross-aligned with the principal axes of each Gaussian component. This “cross scheme” resembles classical sigma–point constructions (Julier and Uhlmann, 2004) in that points lie along eigenvector directions of the covariance; however, it fundamentally differs by (i) allowing an arbitrary number of points per axis, and (ii) assigning exact probabilities via the Voronoi partition w.r.t. to the points, rather than choosing weights to satisfy a finite set of moment constraints.
Figure 10: Construction of the cross-shaped signature of a 2D Gaussian (black dots) using the signature of a standard Gaussian (blue dots) as a reference. In the whitened eigenbasis space defined by T{{T}} (orange axes), we place symmetric points along each principal axis by mirroring the template locations (green markers) across all axes to form the cross. Mapping these points back via ${{T}}^{-1}$ yields the signature locations in the original space. The blue (left) and black (right) lines indicate the boundaries of the Voronoi regions induced by the signature locations in the transformed and original spaces, respectively.
The procedure to compute the component-wise signature of a Gaussian mixture (see Eqn (14)) is illustrated in Figure 10 and summarized in Algorithm 4. First (line 2) we obtain, via Algorithm 5, the radial template (axis radii) in the transformed space (green stripes in Figure 10). Note that we set the number of radial levels $L=\max\{1,\lfloor N_{\text{cross}}/(2n)\rfloor\}$ (line 1), allocating $2n$ points per level (two per axis) while ensuring at least one level, where $\lfloor\cdot\rfloor$ denotes the floor function. Next we compute the probabilities of the Voronoi regions induced by the eventual signature points by: (i) evaluating cumulative probabilities of the $n$-dimensional standard Gaussian ball at the radial Voronoi boundaries using the $\chi^{2}_{n}$ CDF (line 4); and (ii) differencing these to obtain shell masses $\Delta_{\ell}$ (line 5). The weight of each point at radial level $\ell$ is then $w_{\ell}=\Delta_{\ell}/(2n)$ (line 6). For each mixture component (lines 7-9), at every radial level $\ell$ and along every principal axis $j$, we place two symmetric points ${m}_{i}\pm r_{\ell}s_{i,j}{{a}}_{i,j}$ with weight $w_{\ell}$ and map them back to the original space. The heuristic in Algorithm 5 ensures that the resulting signature matches the second moment of the corresponding Gaussian component.
input : $q=\sum_{i=1}^{M}{\tilde{{{\pi}}}^{{(i)}}}\,\mathcal{N}({m}_{i},\Sigma_{i})$, desired number of signature locations $N_{\text{cross}}$
output : Signature $d=\sum_{j=1}^{N}{{{\pi}}^{(j)}}\delta_{{{c}}_{j}}$
begin
 1 Set number of radial levels $L=\max\{1,\lfloor N_{\text{cross}}/(2n)\rfloor\}$
 2 Obtain radii $\{r_{\ell}\}_{\ell=1}^{L}$ via Algorithm 5
 3 Define radial Voronoi boundaries $u_{\ell}=\tfrac{1}{2}(r_{\ell}+r_{\ell+1})$ for $\ell=1,\dots,L-1$
 4 Compute cumulative $n$-D standard normal ball probabilities $B(u_{\ell})=P(\|Z\|_{2}\leq u_{\ell}),\;Z\sim\mathcal{N}(0,{{I}}_{n})$ using $B(u_{\ell})=\Phi_{\chi^{2}_{n}}(u_{\ell}^{2})$
 5 Set shell masses $\Delta_{1}=B(u_{1})$, $\Delta_{\ell}=B(u_{\ell})-B(u_{\ell-1})$ for $\ell=2,\dots,L-1$, and $\Delta_{L}=1-B(u_{L-1})$
 6 Compute per-level point weights $w_{\ell}=\Delta_{\ell}/(2n)$
 for $i\in\{1,\dots,M\}$ do
    8 Set unit axis directions ${{a}}_{i,j}={{V}}_{i}{{e}}_{j}$ and scales $s_{i,j}=\sqrt{{{{\lambda}}_{i}^{{(j)}}}}$ for $j=1,\dots,n$
    9 Construct $d_{i}=\sum_{\ell=1}^{L}\sum_{j=1}^{n}w_{\ell}\Big{(}\delta_{{m}_{i}+r_{\ell}s_{i,j}{{a}}_{i,j}}+\delta_{{m}_{i}-r_{\ell}s_{i,j}{{a}}_{i,j}}\Big{)}$
 10Return $d=\sum_{i=1}^{M}{\tilde{{{\pi}}}^{{(i)}}}d_{i}$
Algorithm 4 Cross Signatures of Gaussian mixtures
input : Number of radial levels $L$, dimension $n$
output : Axis radii $\{r_{\ell}\}_{\ell=1}^{L}$
begin
 1 Define probability edges $\{p_{\ell}\}_{\ell=0}^{L}$ with $p_{\ell}=0.5+\tfrac{\ell}{2L}$ (uniform partition of $[0.5,1]$)
 2 Compute quantiles $q_{\ell}=\Phi^{-1}(p_{\ell})$ for $\ell=0,\dots,L$
 3 Compute shell probabilities $s_{\ell}=\Phi(q_{\ell})-\Phi(q_{\ell-1})$ for $\ell=1,\dots,L$
 4 Set radii $r_{\ell}=\big{(}\phi(q_{\ell-1})-\phi(q_{\ell})\big{)}/s_{\ell}\;\sqrt{n}$ for $\ell=1,\dots,L$ (positive since $\phi(q_{\ell-1})>\phi(q_{\ell})$)
 5 Return $\{r_{\ell}\}_{\ell=1}^{L}$.
Algorithm 5 Axis radii for the cross scheme
Appendix E Compression for Dropout-Induced Bernoulli Mixtures
We present an alternative to Algorithm 3 for performing compression in the case of mixtures of multivariate Bernoulli distributions. We begin by showing that such mixtures naturally arise in dropout networks and that compression is necessary to enable tractable analysis of their output distributions. We then introduce a structure-aware compression method that yields a closed-form upper bound on the resulting Wasserstein distance.
Let $\phi_{{b}}:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n}$ be the masking function defined by $\phi_{{b}}({{z}})={{b}}\odot{{z}}$, where ${{b}}\in\{0,1\}^{n}$ is a binary vector and $\odot$ denotes the element-wise multiplication operation. In dropout, b{{b}} is sampled from a multivariate Bernoulli distribution $d_{{b}}$, given by
$d_{{b}}=\sum_{{{b}}\in\{0,1\}^{n}}\theta^{\sum_{l=1}^{n}{{{b}}^{(l)}}}(1-\theta)^{n-\sum_{l=1}^{n}{{{b}}^{(l)}}}\delta_{{b}},$ (42)
where $\theta\in[0,1]$ is the dropout rate. Note that $d_{b}$ is a discrete distribution with support $\{0,1\}^{n}$. Given any input distribution $p_{in}$, the output distribution of a dropout layer is defined as
$p_{out}={\mathbb{E}}_{{{b}}\sim d_{{b}}}[\phi_{{b}}\#p_{in}],$ (43)
which is a mixture distribution of $2^{n}$ components. If the input distribution $p_{in}$ is a discrete distribution of size $N$, such as the distribution at an intermediate layer of a dropout network, or a signature approximation in a VI-trained network, then the output becomes a discrete distribution with support size $N\cdot 2^{n}$. Specifically, for $d_{in}=\sum_{i=1}^{N}{{{\pi}}^{(i)}}\delta_{{{c}}_{i}}$, we have
$d_{out}={\mathbb{E}}_{{{b}}\sim d_{{b}}}[\phi_{{b}}\#d_{in}]=\sum_{i=1}^{N}{{{\pi}}^{(i)}}\sum_{{{b}}\in\{0,1\}^{n}}\theta^{\sum_{l=1}^{n}{{{b}}^{(l)}}}(1-\theta)^{n-\sum_{l=1}^{n}{{{b}}^{(l)}}}\delta_{{{b}}\odot{{c}}_{i}}.$ (44)
Analyzing or storing $d_{out}$ becomes quickly intractable for moderately large $n$ due to its exponential growth in support size. Therefore, we require an efficient compression procedure to approximate $p_{out}$ with a tractable distribution.
Given a discrete input distribution $d_{in}$, one could apply the approach from Algorithm 3 to compress $d_{out}$ into a discrete distribution $\bar{d}_{out}\in{\mathcal{D}}_{M}({\mathbb{R}}^{n})$, where $N\leq M<N\cdot 2^{n}$. This would involve clustering the support of $d_{out}$ into $M$ groups using M-means, applying moment matching within each cluster, and solving a discrete optimal transport problem to compute the 2-Wasserstein error. However, this approach is computationally demanding when the number of components is large.
Instead, we exploit the structure of the Bernoulli mixture to design a more efficient compression strategy that admits a closed-form upper bound on the 2-Wasserstein distance. Specifically, given $d_{in}$, we select the $\frac{M-N}{2^{n}-1}$ components ${{c}}_{i}$ with the largest norms and collect their indices into a set $\mathcal{I}$} using a simple sorting algorithm. We then define the compressed distribution as:
$\bar{d}_{out}=\sum_{i\in{\mathcal{I}}^{c}}{{{\pi}}^{(i)}}\delta_{{{c}}_{i}}+\sum_{i\in{\mathcal{I}}}{{{\pi}}^{(i)}}\sum_{{{b}}\in\{0,1\}^{n}}\theta^{\sum_{l=1}^{n}{{{b}}^{(l)}}}(1-\theta)^{n-\sum_{l=1}^{n}{{{b}}^{(l)}}}\delta_{{{b}}\odot{{c}}_{i}}$ (45)
where ${\mathcal{I}}^{c}=\{1,\ldots,N\}\setminus{\mathcal{I}}$. That is, we apply dropout only to the inputs in $\mathcal{I}$}, and leave the others unperturbed. The resulting distribution $\bar{d}_{out}$ has support size
$|{\mathcal{I}}^{c}|+|{\mathcal{I}}|\cdot 2^{n}=\left(N-\frac{M-N}{2^{n}-1}\right)+\frac{M-N}{2^{n}-1}\cdot 2^{n}=M.$
The following proposition provides a closed-form upper bound on the 2-Wasserstein distance between $d_{out}$ and $\bar{d}_{out}$.